### v1

# Counting Collatz Numbers

Abstract
The counting function $\pi(x)$ for the numbers satisfying the Collatz conjecture is studied. A related exponential congruence equation is investigated, yielding a method to construct its solutions from free tuples, and enabling us to obtain the inequality
$\pi(x)\geq x^{0.946},\ x\rightarrow+\infty.$
The historical record is 0.84.
Key words: Collatz conjecture, 3x+1 problem, exponential congruence equation, ordered partition.
1 INTRODUCTION
We begin with the following definitions.
Definition 1.1 (Syracuse map)
The Syracuse map $S$ is an operator on the set of odd numbers acting according ro the formula
$S(n)=(3n+1)2^{-{\rm ord}_{2}(3n+1)}.$
The $l$-th iteration of $S$ is denoted as $S^{l}$.
Definition 1.2 (Collatz number)
An odd number $n$ is called a Collatz number if $S^{l}(n)=1$ for some $l$. The level of a Collatz number is the smallest number $l$ such that $S^{l}(n)=1$.
We now recall the Collatz conjecture.
Conjecture 1.3 (Collatz conjecture)
Every odd number is a Collatz number.
Definition 1.4 (Collatz number counter)
Let
$\pi(x)=\#\{n\leq x\mid n\text{ is a Collatz number}\},x>0.$
In 1978, Crandall[Cr78] proved that there is a positive constant $\beta>0$ such that
$\pi(x)\geq x^{\beta},\ x\rightarrow+\infty.$
In 1989, Krasilov [Kr89] proved that
$\pi(x)\geq x^{0.43},\ x\rightarrow+\infty.$
In 1993, Wirsching [Wi93] proved that
$\pi(x)\geq x^{0.48},\ x\rightarrow+\infty.$
In 1995, Applegae-Lagarias [AL2] proved that
$\pi(x)\geq x^{0.81},\ x\rightarrow+\infty.$
In 2003, Krasilov-Lagarias [KL03] proved that
$\pi(x)\geq x^{0.84},\ x\rightarrow+\infty.$
In 2022, Tao [Tao22] studied the numbers whose orbit under the Syracuse attains almost bounded values. In this paper, we shall prove the following result.
Theorem 1.5
There is a positive constant $c$ such that
$\pi(x)\geq cx^{H_{2}(1/\log_{2}3)},\ x\rightarrow+\infty,$
where $H_{2}(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$. In particular,
$\pi(x)\geq x^{0.946},\ x\rightarrow+\infty.$
In the next section, we shall investigate an exponential congruence equation, and give a method to construct its solutions from free tuples. In the section after the next, we shall relate Collatz numbers to the mentioned congruence equation. In the final section we shall give a proof of the main result of this paper.
2 A Congruence Equation
In this section we investigate an exponential congruence equation. We give a method to construct its solutions from free tuples.
Definition 2.1 ($(2,3)$-primary Congruence Equation)
Let $x_{1},\cdots,x_{l}$ be $l$ variables taking values in numbers. Then the congruence equation
$2^{x_{1}+\cdots+x_{l}}\equiv\sum_{j=0}^{l-1}3^{j}2^{x_{j+2}+\cdots+x_{l}}({\rm mod}3^{l})$
is called the $(2,3)$-primary congruence equation of level $l$.
Lemma 2.2
Let $(u_{1},\cdots,u_{l})$ be an $l$-tuple of numbers. Then there is a unique solution $(v_{1},\cdots,v_{l})$ of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
Proof. Firstly, if $l=1$, then it is easy to see that $v_{1}=2u_{1}$ is the unique solution of the $(2,3)$-primary level-11 congruence equation
$2^{x_{1}}\equiv 1({\rm mod}3)$
such that
$\frac{v_{1}+1}{2}]=u_{1}.$
Secondly, we assume that $l>1$ and $(v_{1},\cdots,v_{l-1})$ is the unique solution of the $(2,3)$-primary level-(l1)(l-1) congruence equation
$2^{x_{1}+\cdots+x_{l-1}}\equiv\sum_{j=0}^{l-2}3^{j}2^{x_{j+2}+\cdots+x_{l-1}}({\rm mod}3^{l-1})$
such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l-1.$
Then
$2^{v_{1}+\cdots+v_{l-1}}\equiv\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}({\rm mod}3^{l-1}).$
Let
$a=3^{1-l}(2^{v_{1}+\cdots+v_{l-1}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}).$
Then $a\neq 0({\rm mod}3)$, and there is a unique binary digit $r_{l}$ such that
$2^{r_{l}}\equiv a({\rm mod}3).$
Let
$v_{l}=2u_{l}-r_{l}.$
Then
$2^{v_{l}}a\equiv 1({\rm mod}3).$
Let
$b=\frac{2^{v_{l}}a-1}{3}.$
Then
$2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l}3^{j}2^{v_{j+2}+\cdots+v_{l}}=3^{l}b\equiv 0({\rm mod}3^{l}),$
and
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
That is, $(v_{1},\cdots,v_{l})$ a solution of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
From the inductive construction process of $(v_{1},\cdots,v_{l})$, we see that $(v_{1},\cdots,v_{l})$ is uniquely determined by $(u_{1},\cdots,u_{l})$. The lemma is proved.
3 Collatz numbers
In this section relate Collatz numbers to solutions of the $(2,3)$-primary congruence equation.
Definition 3.1
Let $(v_{1},\cdots,v_{l})$ be a solution of the $(2,3)$-primary congruence equation of level $l$. We write
$B(v_{1},\cdots,v_{l}))=3^{-l}(2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l}}).$
Lemma 3.2
Let $n>1$ be a Collatz number of level $l$, and
$v_{m}={\rm ord}_{2}(3S^{l-m}(n)+1),\ m=1,\cdots,l.$
Then $v_{1}>2$, and
$n=3^{-l}(2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l}}).$
In particular, $(v_{1},v_{2},\cdots,v_{l})$ is a solution of the $(2,3)$-primary congruence equation of level $l$ such that
$B(v_{1},\cdots,v_{l}))=n.$
Proof. Firstly, if $l=1$, then
$v_{1}={\rm ord}_{2}(3n+1).$
Thus
$1=S(n)=(3n+1)2^{-v_{1}}.$
Hence $v_{1}>2$, and
$n=3^{-1}(2^{v_{1}}-1).$
We now assume that $l>1$. Then
$v_{m}={\rm ord}_{2}(3S^{l-1-m}(S(n))+1),\ m=1,\cdots,l-1.$
By induction, $v_{1}>2$, and
$S(n)=3^{-l+1}(2^{v_{1}+\cdots+v_{l-1}}-\sum_{j=0}^{l-2}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}).$
Hence
$n=\frac{2^{v_{l}}S(n)-1}{3}=3^{-l}(2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l}}).$
The lemma is proved.
Lemma 3.3
If $(v_{1},v_{2},\cdots,v_{l})$ with $v_{1}>2$ is a solution of the $(2,3)$-primary congruence equation of level $l$, and
$n=B(v_{1},\cdots,v_{l})),$
then $n$ is a Collatz number of level $l$, and
$v_{l-m}={\rm ord}_{2}(3S^{m}(n)+1),\ m=0,\cdots,l-1.$
By this lemma, different tuples give rise to different Collatz numbers. Proof. Firstly, if $l=1$, then
$n=3^{-1}(2^{v_{1}}-1)\neq 1.$
Thus
$v_{1}={\rm ord}_{2}(3n+1),$
and
$S(n)=(3n+1)2^{-v_{1}}=1.$
Therefore the assertion in the lemma is true when $l=1$. We now assume that $l>1$. It is easy to see that
${\rm ord}_{2}(3n+1)=v_{l},$
and
$S(n)=(3n+1)2^{-v_{l}}=3^{-l+1}(2^{v_{1}+\cdots+v_{l-1}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}\cdots+v_{l-1}}).$
In particular, $(v_{1},v_{2},\cdots,v_{l-1})$ is a solution of the $(2,3)$-primary congruence equation of level $l$. By induction, $S(n)$ is a Collatz number of level (l1)(l-1), and
$v_{l-1-m}={\rm ord}_{2}(3S^{m}(S(n))+1),\ m=0,\cdots,l-2.$
and Therefore, $n$ is a Collatz number of level $l$, and
$v_{l-m}={\rm ord}_{2}(3S^{m}(n)+1),\ m=0,\cdots,l-1.$
The lemma is proved.
4 Bounds
In this section we prove the main theorem of this paper.
Definition 4.1 (Fixed level Collatz number counter)
For $l\in\mathbb{N}$, let
$\pi(x,l)=\#\{n\leq x\mid n\text{ is a Collatz number of level }l\},x>0.$
Definition 4.2 (Ordered partition counter)
For $l\in\mathbb{N},y>0$, let
$\omega(y,l)=\#\{(u_{1},\cdots,u_{l})\mid\sum_{j=1}^{l}u_{j}\leq y,(u_{1},\cdots,u_{l})\text{ is a tuple of numbers with }u_{1}>1\}.$
Lemma 4.3 (Partition Bound)
Let $x>0$. Then
$\pi(x,l)\geq\omega(\log_{4}x+l\log_{4}3,l).$
Proof. In fact, let $(u_{1},\cdots,u_{l})$ be an $l$-tuple of numbers with $u_{1}>1$ such that
$\sum_{j=1}^{l}u_{j}\leq\log_{4}x+l\log_{4}3.$
That is, $(u_{1},\cdots,u_{l})$ is an ordered $l$-partition with $u_{1}>1$ of a number $\leq\log_{4}x+l\log_{4}3.$ Let $(v_{1},\cdots,v_{l})$ be the unique solution of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
Then $B(v_{1},\cdots,v_{l})$ is a Collatz number of level $l$. It is easy to see that
$B(v_{1},\cdots,v_{l})\leq 3^{-l}2^{v_{1}+\cdots+v_{l}}\leq 3^{-l}4^{u_{1}+\cdots+u_{l}}\leq x.$
It is also easy to see that different ordered partitions give rise to different Collatz numbers. It follows that
$\pi(x,l)\geq\omega(\log_{4}x+l\log_{4}3,l).$
The lemma is proved.
Lemma 4.4
Let $l=[\frac{\log_{2}x}{\log_{2}3}]$, and $n=[\log_{4}x+l\log_{4}3]-1$. There is a positive constant $c_{0}$ independent of $x$ such that, if $x$ is sufficiently large, then
$\log_{2}{n\choose l}\geq H_{2}(1/\log_{2}3)\log_{2}x-O(1),$
where $H_{2}(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
Proof. Notice that
$n-1=\log_{2}x+O(1),\text{ and }l/n=1/\log_{2}3+O(1/\log_{2}x).$
Thus, by Stirling’s formula,
$\log_{2}{n\choose l} \displaystyle\geq$ $nH_{2}(l/n)-O(1)$
$\geq$ $(\log_{2}x)H_{2}(1/\log_{2}3)-O(1).$
The lemma now follows.
Proof of Theorem 1.5. Let $l=[\frac{\log_{2}x}{\log_{2}3}]$, and $n=[\log_{4}x+l\log_{4}3]-1$. Then
$\pi(x)$ $\geq$ $\pi(x,l).$
$\geq$ $\omega(\log_{4}x+l\log_{4}3,l)$
$\geq$ ${n\choose l}$
$\geq$ $2^{-c_{0}}x^{H_{2}(1/\log_{2}3)}.$
Choosing $c=2^{-c_{0}}$, we complete the proof of Theorem 1.5.